# Can LLMs Really Learn to Translate a Low-Resource Language from One Grammar Book?

### v1

@@ 4.1 Data
We use the preprocessed Kalamang (kgv) grammar book (Visser, 2022; Tanzer et al., 2024), with additional processing of irregularities (particularly for glossing) introduced in LaTeX conversion. We similarly preprocess a grammar text in Nepali (npi) (Bal, 2004). We prompt with the entire grammar, , in-context. We heuristically filter parallel glossed examples and bilingual word/phrase pairs from the book into a parallel subset, . The remainder of the book contains grammatical explanations without parallel signals, labelled . Subset statistics are shown in Table 1. We preprocess kgv parallel examples from the grammar book into an unsegmented, structured format, giving (used for 5*-shot examples and in full as a prompt) and which includes glosses (1239 examples) – for excerpts of prompt types, see Appendix C. We additionally test prompts with (400 examples) and Wordlist (W) (3813 examples). Additionally, we sample 500 examples from Dictionaria333https://dictionaria.clld.org/contributions/kalamang (Visser, 2020) as the development set for fine-tuning. In total there are 3.3k engkgv parallel examples444Data and code will be shared after review.; we focus on the 1.2k from for fair comparison with Book settings. For testing, we use our combined 100 example test set for Kalamang, and the Flores devtest for Nepali (1012 examples) (Guzmán et al., 2019), with few-shot examples from Flores dev.
For IGT prediction, we preprocess 1221 examples from the Grammar Book with complete glosses for the training (5623 words) and development (612 words) set (split 90:10% by sentences). Following Ginn et al. (2023), we introduce a test set of 97 glossed examples (447 words) from a different source, Dictionaria, which were sanity-checked manually for correct alignment.
@@ 4.2 Models
In our experiments we use the API-only Gemini-1.5-Flash-001 (henceforth Gemini) (Gemini Team et al., 2024). We justify this choice due to Gemini’s context window of 1M tokens, significantly larger than other models, which can handle the entire grammar book, and use Flash over Pro due to the prohibitive cost differences. We also use the open-weight Llama-3.1-8B555https://huggingface.co/meta-llama/Meta-Llama-3.1-8B (Dubey et al., 2024) and its instruction-tuned variant, with a context of 128k tokens. This is insufficient for kgv , but fits and , plus the npi . Since this work is intended to help under-resourced language speakers, the minimisation of costs by using smaller, open-weight models is well-motivated. We test Llama-Instruct (Llama-I), and fine-tune Llama base with LoRA (Hu et al., 2021) on Pbook (Llama-ftp) with prompt masking for 5 epochs with a constant learning rate of 1e-4, batch size 4, and LoRA , , targeting all linear projections. For our NMT baseline, we fine-tune NLLB-1.3B-Distilled (NLLB) (Costa-jussà et al., 2024) on Pbook. For Kalamang grammaticality judgment and IGT prediction, we use the same Gemini model as above.
@@ 4.3 Evaluation
We evaluate translation automatically with ChrF++ (Popović, 2017). We favour ChrF++ over ChrF, used in Tanzer et al. (2024), since it takes into account word order, which is central to correct grammar. We report scores for trimmed responses after the first newline character to distinguish translation quality from overgeneration and chat explanations (Aycock & Bawden, 2024) and use a forceful prompt (detailed in Appendix C) to ensure the translation is produced on the first line.
@@ 4.4 Baselines
For translation experiments, we test the following baselines: 0-shot translation with a standard translation prompt; word-for-word translation with fuzzy dictionary lookup (W4W); 5 retrieved examples per word (5*-shot) based on longest common subsequences following Tanzer et al. (2024); prompting with the full dictionary (Wordlist, W); the processed set of book parallel examples () and with glosses (); and processed training set examples ().
For IGT prediction, we use a baseline frequency-based classifier (Top-Class), a fine-tuned RoBERTa token classifier (Ginn et al., 2023) (SMP-Base); a hard-attention glossing model (TüCL-Morph) (Girrbach, 2023); and ByT5-ft and Glosslm-ft models (Ginn et al., 2024b) fine-tuned on our kgv IGT training and development sets. We provide segmented input, and English translations to models which accept them.
@@ 4.5 Experiments
Our central research question investigates the contributions of grammatical explanations and parallel data to translation performance. We therefore prompt models with and the complementary subsets and . We test our typological feature prompt, Typ, as a replacement for . For npi, we repeat the book settings as above. We fine-tune translation models with the preprocessed Pbook parallel data for comparison with settings. For grammaticality judgment and IGT prediction tasks, we similarly test Gemini with the kgv and Typ prompts.
@@ 5.1 Results
We disentangle the signal from grammar books’ explanations and parallel sentences for translation. Our kgv results in Table 5 show that most or all performance improvements stem from the book’s parallel sentences, with quality plummeting when parallel data is removed. With Gemini into eng, marginally outperforms , and beats by 7 ChrF++, while into kgv, outperforms by over 8 points, and performs 3 points better than . However, we show statistically in Section 5.2 that this small improvement is modelled directly by an increase in test set vocabulary coverage, rather than from the grammatical explanations. Additionally, this gap closes with the prompt, which preprocesses and structures the parallel data in into kgv–gloss–eng triples. performs particularly well for Llama-I, with over 10 points improvement over into kgv. Due to context limits restricting kgv tests, conclusions with Llama-I are limited, but we find again that performance lags far behind . We note that baselines including 0-shot show kgv translation is non-trivial. We also find that additional parallel data further improves translation quality, and note 5*-shot is generally competitive despite its short average prompt, achieving the best into kgv with Gemini. Thus for kgv translation, both LLMs on test mainly learn from the book’s parallel sentences, failing to exploit the grammatical explanations.
We observe a similarly strong trend for npi, a seen low-resource language with a high-quality Flores test set, in Table 5.1. settings either match or outperform for both models and directions. We note few settings beat 0-shot and differences between Gemini settings are less pronounced than for kgv; perhaps the model’s prior competence and a shorter grammar book mean there is less to be gained. However, analysing Book settings in isolation shows that both and have a detrimental effect of up to 4 points below 0-shot with Llama-I, while has a neutral or small positive impact into and from npi respectively. Finally 5*-shot is again effective especially for Llama-I into npi, likely due to the greater vocabulary coverage of the example set. These results generalise our findings for kgv to seen low-resource languages: we find no evidence that LLMs can effectively exploit grammatical explanations for translation.
@lcccc@
@@
Setting↓ eng–npi npi–eng
Model→ Gemini L-I Gemini L-I
0-shot 42.5 28.6 65.2 51.1
5*-shot 43.2 37.6 64.9 57.3
42.6 24.3 64.4 48.9
42.5 28.6 64.9 52.6
41.8 24.5 64.5 48.4
Typ 0-shot 42.4 23.2 64.6 49.5
Typ + 41.8 22.0 64.9 49.1
@@
We test a standard MT approach for adding a new language by fine-tuning NLLB, a small MT model, on the book’s parallel data, shown in Table 5.1. NLLB achieves competitive or improved performance compared to prompting Gemini with the same preprocessed parallel data, . We also test backtranslation (BT), a standard method to boost performance in MT (Sennrich et al., 2016). A single BT iteration with has a negative impact into kgv, likely due to the poor quality of the initial model introducing excessive noise. However we see a boost of 3 ChrF++ into eng, we expect because of the strong English language modelling of NLLB. Further, adding a small 400 example parallel training set sees large gains of 4-8 points. These results suggest the MTOB benchmark can be adequately addressed as a standard XLR MT problem with simple data preprocessing, a small pre-trained model, and fine-tuning on a single GPU for 1 hour.
@@
We also fine-tune Llama base on to give Llama-ft, with results in Table 5. We find all Llama-ft settings beat equivalent Llama-I tests with data, except for settings with glosses which marginally outperform Llama-ft 0-shot results. Prompting Llama-ft with parallel data in-context further improves performance over 0-shot by up to 10 points. We additionally fine-tune Gemini on , with results in Appendix F, finding Gemini-ft underperforms NLLB and Gemini with the same data in-context by 6-12 ChrF++; we expect this is because it is already extensively instruction-tuned. Thus fine-tuning—particularly of small MT models—is a cheap method for achieving competitive results with prompting instruction-tuned long-context LLMs, given the same parallel data.
@@
Given the limited contribution of grammatical explanations to translation performance, we introduce a novel prompting method summarising languages’ typological features. This prompt is intended to replace , thus we are primarily focused on results when combined with data. Our results for eng–kgv translation in Table 5 show expectedly poor 0-shot performance due to the lack of any Kalamang text. Into kgv, our prompt beats but not ; however into eng, our prompt with achieves the best translation results for settings with book parallel data. Similarly for Nepali in Table 5.1, Typ + is less effective than into npi, and marginally outperforms it into eng up to 0.5 ChrF++, though alone performs best. The performance of typological prompting for translation is therefore inconsistent, supporting the above finding that LLMs fail to effectively exploit grammatical information for MT.
@@ 5.2 Analysis
We investigate whether any added performance from grammatical explanations is statistically significant or can instead be attributed to greater test set type coverage in the prompt. We distinguish between types, meaning unique words in a vocabulary, and tokens, i.e. individual occurrences of types in a text. We fit univariate least squares regression models to ChrF++ scores with test set type coverage as the independent variable, for both directions, shown in Figure 2. No setting falls outside the confidence interval, and the models are significant in both directions ()666For details of these and following statistical tests, see Appendix G.; the correlations are also significant (). Thus maximising target vocabulary coverage (via parallel sentences) in-context is the most efficient method for improving LLM-based XLR translation. These linear regressions show that translation performance can be directly modelled by test set vocabulary coverage, and that the book’s grammar explanations provide no significant advantage over its parallel sentences. See Appendix D for full statistics on our prompts’ test set type coverage.
We then explore whether the improved translation scores can be attributed to a longer (or shorter) prompt, by testing for a relationship between prompts’ total tokens and translation quality in terms of ChrF++ for with Gemini. The resulting linear models are not significant in either direction (, into and from kgv), with no correlations. The grammar book is therefore both a token-inefficient way to learn (with similar performance despite nearly 5x more tokens than kgv ), and a cost-inefficient dataset to generate, compared to using its parallel sentences. The needle-in-a-haystack problem could partially explain this: with increasing context, retrieval of relevant information (i.e. similar parallel examples) becomes harder (Hsieh et al., 2024), so while is a subset of , there is a greater ratio of relevant to irrelevant information in the prompt—assuming grammatical explanations cannot be effectively exploited for translation.
@@ Discussion
We note that our results do not indicate LLMs cannot understand books in general; rather, we find no quantitative evidence that the results here and in MTOB show LLMs can effectively exploit grammar books (or linguistic knowledge) for translation. Indeed, we show that LLMs can exploit grammatical information in the form of typology for more relevant, linguistically-focused tasks. More broadly, from an educational perspective, translation is a problem-solving task aiming to reach a goal state (translation) via a series of actions given an initial state (source) and optionally rules on applying actions. Humans tend to learn more efficiently via worked-examples (van Gog et al., 2019), i.e. with explicit explanations, rather than pure discovery learning, meaning without explicit guidance (Mayer, 2004). Our results however indicate that for translation, LLMs learn more effectively from unannotated parallel examples (i.e. discovery) than from grammar principles with explained examples (i.e. example-based). Our results thus tentatively support a divergence between learning strategies for translation between human learners and LLMs learning in-context.
@@ 6 Conclusion
We find no evidence that LLMs can effectively exploit grammatical explanations for XLR MT in Kalamang and Nepali, relying instead on the parallel sentences within the book. This runs counter to the claim of prior work including MTOB which use grammar books to enable XLR tasks. We show that fine-tuning small MT models matches the performance of costly long-context LLMs. Further, we show statistically that grammar adds no significant advantage above the increased type coverage it provides, and that grammar books are less token-efficient for prompting than parallel sentences. However, LLMs can exploit grammatical information, given an appropriate task—e.g. grammaticality judgment or IGT prediction—and more useful grammatical data in the form of our typological prompt, which achieves leading results on these linguistic tasks. We therefore emphasise the importance of task-appropriate data: parallel data for MT, and grammatical, preferably typological, knowledge for linguistic tasks. Moreover, we suggest data collection efforts for multilingual XLR tasks, at least for MT, are better focused on parallel data over linguistic description, which enables less costly, more token-efficient translation.
@@
We emphasise that this work does not aim to and does not solve social problems, and instead investigates the utility of grammar books as resources for XRL NLP. We operate on the assumption of continued consent of the Kalamang community to use their language in our research, as discussed in Tanzer et al. (2024). We discuss the utility of our work (specifically the IGT prediction task) for linguists and L1 speakers, in its capacity to improve language documentation processes.

### v2

@@ 4.1 Data
We use the preprocessed Kalamang (kgv) grammar book (Visser, 2022; Tanzer et al., 2024), with additional processing of irregularities (particularly for glossing) introduced in LaTeX conversion. We similarly preprocess a grammar text in Nepali (npi) (Bal, 2004) and Paraguayan Guarani (gug) (Estigarribia, 2020). We prompt with the entire grammar, , in-context (where the subscript indicates the data subset). Following Nordhoff & Krämer (2022), we extract parallel glossed examples and bilingual word/phrase pairs from the book based on text formatting into a parallel subset, . The remainder of the book contains grammatical explanations without parallel examples, labelled . Subset statistics are shown in Table 1. We preprocess kgv–eng parallel examples from the grammar book into an unsegmented parallel data format, giving (used for 5*-shot examples and in full as a prompt) and which includes glosses (1239 examples) – for excerpts of prompt types, see Appendix E. We additionally test prompts with (400 examples) and Wordlist (W) (3813 examples). Additionally, we sample 500 examples from Dictionaria222https://dictionaria.clld.org/contributions/kalamang (Visser, 2020) as the development set for fine-tuning. In total there are 3.3k engkgv parallel examples333Data (including grammar book splits) and code are made available at this link.; we focus on the 1.2k in for fair comparison with Book settings. For testing, we use our combined 100 example test set for kgv, and Flores devtest for npi and gug (1012 examples) (Guzmán et al., 2019), with few-shot examples from Flores dev. For IGT prediction, we preprocess 1221 examples from the grammar book with glosses for training (5623 words) and development (612 words) sets (split 90:10% by sentences). Following Ginn et al. (2023), we introduce a test set of 97 glossed examples (447 words) from a different source, Dictionaria, which were manually inspected for correct alignment.
@@ 4.2 Models
In our experiments we use the API-only Gemini-1.5-Flash-001 (henceforth Gemini) (Gemini Team et al., 2024). We justify this choice due to Gemini’s context window of 1M tokens, significantly larger than other models, which can handle the entire grammar book, and use Flash over Pro due to prohibitive cost differences. We also use the smaller, open-weight Llama-3.1-8B base and instruction-tuned models (Dubey et al., 2024), with a context of 128k tokens. This is insufficient for kgv and gug , but fits and , plus the npi . We test Llama-Instruct (Llama-I), and fine-tune Llama base with LoRA (Hu et al., 2021) on (Llama-ft) with prompt masking for 5 epochs with a constant learning rate of 1e-4, batch size 4, and LoRA , , targeting all linear projections. For our NMT baseline, we fine-tune NLLB-1.3B-Distilled (NLLB) (Costa-jussà et al., 2024) on . For kgv grammaticality judgment and IGT prediction, we use the same Gemini model as above.
@@ 4.3 Evaluation
We evaluate translation automatically with ChrF++ (Popović, 2017). We favour ChrF++ over ChrF, used in Tanzer et al. (2024), since it takes into account word order as well as character -gram overlap. We report scores for trimmed responses after the first newline character to distinguish translation quality from overgeneration and chat explanations (Aycock & Bawden, 2024) and use a forceful prompt (detailed in Appendix E) to ensure the translation is produced on the first line.
@@ 4.4 Baselines
For translation experiments, we test several baselines: 0-shot translation with a standard translation prompt; word-for-word translation with fuzzy dictionary lookup (W4W); 5 retrieved examples per word (5*-shot) based on longest common subsequences following Tanzer et al. (2024); prompting with the full Wordlist (W), parallel examples, , parallel examples with glosses, , and processed training set examples, . For IGT prediction, we use a baseline frequency-based classifier (Top-Class), a fine-tuned RoBERTa token classifier (Ginn et al., 2023) (SMP-Base); a hard-attention glossing model (TüCL-Morph) (Girrbach, 2023); and ByT5-ft and Glosslm-ft models (Ginn et al., 2024b) fine-tuned on our kgv IGT training and development sets. We provide segmented input, and English translations to models which accept them.
@@ 4.5 Experiments
Our central research question investigates the contributions of grammatical explanations and parallel data to translation performance. We therefore prompt models with and its filtered subsets. We test our typological feature prompt, Typ, to replace . For npi and gug, we repeat the book settings as above. We fine-tune translation models with the parallel data for comparison with settings. For grammaticality judgment and IGT prediction tasks, we similarly test Gemini with the kgv Book and Typ prompts.
@@ 5 Results & Analysis
We disentangle the signal from grammar books’ explanations and parallel sentences for translation. Our kgv results in Table 5 show that most or all performance improvements stem from the book’s parallel sentences, with quality plummeting when parallel data is removed. With Gemini into eng, marginally outperforms , and beats by 7 ChrF++, while into kgv, outperforms by over 8 points, and performs 3 points better than . However, we show statistically in Section 5.1 that this small improvement is modelled directly by an increase in test set vocabulary coverage, rather than from the grammatical explanations. Additionally, this gap closes with the prompt, which preprocesses and structures the parallel data in into kgv–gloss–eng triples. performs particularly well for Llama-I, with over 10 points improvement over into kgv. Due to context restricting kgv tests, conclusions with Llama-I are limited, but we find again that performance lags far behind . We note that baselines including 0-shot show kgv translation is non-trivial. We also find that additional parallel data further improves translation quality, and note 5*-shot is generally competitive despite its short average prompt, achieving the best score into kgv with Gemini. Thus for kgv translation, both LLMs on test mainly learn from the book’s parallel sentences, failing to exploit the grammatical explanations.
We observe a similarly strong trend for npi and gug, seen low-resource languages with high-quality Flores test sets, in Table 5. settings largely match or outperform for both models and languages (except Llama-I in gug–eng where the model often fails to output translations on the first line for Book settings). Few settings beat 0-shot and differences between Gemini settings (especially npi) are smaller than for kgv; perhaps the model’s prior competence (and a shorter npi grammar book) mean there is less to be gained. However, analysing Book settings in isolation shows that both and have a detrimental effect of up to 7 points below 0-shot, while has a neutral or small positive impact in both npi and gug. Finally 5*-shot is again effective, especially for Llama-I into npi and gug, likely due to the greater vocabulary coverage of the example set. These results generalise our findings for kgv to seen low-resource languages: we find no evidence that LLMs can effectively exploit grammatical explanations for translation.
lcccccccc
@@
Setting↓ eng–npi npi–eng eng–gug gug–eng
Model→ Gemini L-I Gemini L-I Gemini L-I Gemini L-I
0-shot 42.5 28.6 65.2 51.1 26.6 6.1 41.3 23.6
5*-shot 43.2 37.6 64.9 57.3 29.2 13.7 43.1 23.4
42.6 24.3 64.4 48.9 22.2 – 38.7 –
42.5 28.6 64.9 52.6 25.8 6.7 41.8 11.8
41.8 24.5 64.5 48.4 19.3 5.6 34.5 10.1
Typ 0-shot 42.4 23.2 64.6 49.5 21.1 4.3 33.9 23.4
Typ + 41.8 22.0 64.9 49.1 21.9 – 34.5 –
@@
We test a standard MT approach for adding a new language by fine-tuning NLLB, a small MT model, on the book’s parallel data, shown in Table 5. NLLB achieves competitive or improved performance compared to prompting Gemini with the same preprocessed parallel data, . We also test backtranslation (BT), a standard method to boost performance in MT (Sennrich et al., 2016). A single BT iteration with has a negative impact into kgv, likely due to the poor quality of the initial model introducing excessive noise. However we see a boost of 3 ChrF++ into eng, we expect because of the strong English language modelling of NLLB. Further, adding a small 400 example parallel training set sees large gains of 4-8 points. These results suggest the MTOB benchmark can be adequately addressed as a standard XLR MT problem with simple data preprocessing, a small pre-trained model, and fine-tuning on a single GPU for 1 hour.
@@
We also fine-tune Llama base on to give Llama-ft, with results in Table 5. We find all Llama-ft settings beat equivalent Llama-I tests with data, except for settings with glosses which marginally outperform Llama-ft 0-shot results. Prompting Llama-ft with parallel data in-context further improves performance over 0-shot by up to 10 points. We additionally fine-tune Gemini on , with results in Appendix G, finding Gemini-ft underperforms NLLB and Gemini with the same data in-context by 6-12 ChrF++; we expect this is because it is already extensively instruction-tuned. Thus fine-tuning—particularly of small MT models—is a cheap method for achieving competitive results with prompting instruction-tuned long-context LLMs, given the same parallel data.
@@
Given the limited contribution of grammatical explanations to translation performance, we introduce a novel prompting method summarising languages’ typological features. This prompt is intended to replace , thus we are primarily focused on results when combined with data. Our results for eng–kgv translation in Table 5 show expectedly poor 0-shot performance due to the lack of any Kalamang text. Into kgv, our prompt beats but not ; however into eng, our prompt with achieves the best translation results for settings with book parallel data. For npi in Table 5, Typ + is less effective than into npi, and marginally outperforms it into eng up to 0.5 ChrF++, though alone performs best; similarly in gug tests, outperforms Typ + , which beats or matches . The performance of typological prompting for translation is therefore inconsistent, supporting the above finding that LLMs fail to effectively exploit grammatical information for MT.
@@ 5.1 Analysis
We investigate whether any added performance from grammatical explanations is statistically significant or can instead be attributed to greater test set type coverage in the prompt. We distinguish between types, meaning unique words in a vocabulary, and tokens, i.e. individual occurrences of types in a text. We fit univariate least squares regression models to ChrF++ scores with test set type coverage as the independent variable, for both directions, shown in Figure 2. All settings fall within the 95% confidence interval of the regression lines, and the models are significant in both directions (, F-test)444For details of these and following statistical tests, see Appendix B.; the Pearson correlations are also significant (). Thus maximising target vocabulary coverage (via parallel sentences) in-context is the most efficient method for improving LLM-based XLR translation. These linear regressions show that translation performance can be directly modelled by test set vocabulary coverage, and that the book’s grammar explanations provide no significant advantage over its parallel sentences. See Appendix F for full statistics on our prompts’ test set type coverage.
We then explore whether the improved translation scores can be attributed to a longer (or shorter) prompt, by testing for a relationship between prompts’ total tokens and translation quality in terms of ChrF++ for with Gemini. The resulting linear models are not significant in either direction (, into and from kgv, F-test), with no significant Pearson correlations. The grammar book is therefore both a token-inefficient way to learn (with similar performance despite nearly 5x more tokens than kgv ), and a cost-inefficient dataset to generate, compared to using its parallel sentences. The needle-in-a-haystack problem could partially explain this: with increasing context, retrieval of relevant information (i.e. similar parallel examples) becomes harder (Hsieh et al., 2024), so while is a subset of , there is a greater ratio of relevant to irrelevant information in the prompt—assuming grammatical explanations cannot be effectively exploited for translation.
@@ Discussion
We note that our results do not indicate LLMs cannot understand books in general; rather, we find no quantitative evidence that the results here and in MTOB show LLMs can effectively exploit grammar books (or linguistic knowledge) for translation. Indeed, we show that LLMs can exploit grammatical information in the form of typology for more relevant, linguistically-focused tasks. More broadly, from an educational perspective, translation is a problem-solving task aiming to reach a goal state (translation) via a series of actions given an initial state (source) and optionally rules on applying actions. Humans tend to learn this kind of task more efficiently via worked-examples (van Gog et al., 2019), i.e. with explicit explanations, rather than pure discovery learning, meaning without explicit guidance (Mayer, 2004). Our results however indicate that for translation, LLMs learn more effectively from unannotated parallel examples (i.e. discovery) than from grammar principles with explained examples (i.e. example-based). Our results thus tentatively support a divergence between learning strategies for translation between human learners and LLMs learning in-context. We suggest that this may partially stem from prompts with parallel data aligning more closely with LLMs’ instruction-tuning data than grammar book explanations.
@@ 6 Conclusion
We find no evidence that LLMs can effectively exploit grammatical explanations for low and extremely low-resource MT in Kalamang, Nepali, and Guarani, instead finding that LLMs rely on the parallel sentences within the book. This runs counter to the claim of prior work including MTOB which use grammar books to enable LLMs’ performance on XLR tasks. We show that fine-tuning small MT models matches the performance of costly long-context LLMs. Further, we show statistically that grammatical explanations add no significant advantage above the increased type coverage they provide, and that grammar books are less token-efficient for prompting than parallel sentences. However, LLMs can exploit grammatical information, given an appropriate task—e.g. grammaticality judgment or IGT prediction—and more useful grammatical data in the form of our typological prompt, which achieves leading results on these linguistic tasks. We therefore emphasise the importance of task-appropriate data: parallel data for MT, and grammatical, preferably typological, knowledge for linguistic tasks. Moreover, we suggest data collection efforts for multilingual XLR tasks, at least for MT, are better focused on parallel data over linguistic description, which enables less costly, more token-efficient translation.
@@
We emphasise that this work does not aim to address social problems, and instead investigates the empirical utility of grammar books as resources for XLR NLP. We operate on the assumption of continued consent of the Kalamang community to use their language in our research, as discussed in Tanzer et al. (2024). In Sections 2 and 3.5 we discuss the utility of our work for both linguists and L1 speakers, specifically relating to the IGT prediction task in its capacity to improve language documentation processes.
Acknowledgements
This work was funded in part by the UvA’s Language Sciences for Social Good project, the City of Amsterdam, and the Netherlands Organization for Scientific Research (NWO) under project numbers VI.C.192.080 and 2023.017. The authors would like to thank members of the Language Technology Lab for many constructive discussions, particularly Vlad Niculae for detailed feedback on our paper. The authors are grateful for the helpful feedback provided by the anonymous reviewers.